Use a suitable identity to get each of the following products.
 (x + 3) (x + 3)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(x + 3) (x + 3)=(x + 3)^2
= x^2+ 6x + 9
Using (a+b)^2= a^2+ b^2 + 2ab

Thanks